OpenStudy (anonymous):

Volume A box 10 inches high has a square base. The box was made 9 inches longer, and 3 inches narrower (or its width was decreased by 3 inches) while the height remained 10 inches. Give the original length of one side of the box if these changes increased the volume of the box by 50

OpenStudy (anonymous):

put them into an equation of volumes let the square base be x 10 x^2 = 10(x+9)(x-3) + 50

OpenStudy (ddcamp):

Original volume = 10 × y × y New volume = 10 × (y+9) × (y-3) New Volume - Original Volume = 50 10(y+9)(y-3) - 10(y)(y) = 50 Solve for y

OpenStudy (anonymous):

@DDCamp i believe you got that backwards :P 10yy - 10(y+9)(y-3) = 50 Volume increased after change so ..

OpenStudy (ddcamp):

@nphuongsun93 Yes, the volume increased, so the new volume must be larger. So the new volume - the old volume would be a positive 50.

OpenStudy (anonymous):

"The box was made"

OpenStudy (anonymous):

anyway just solve this 10 x^2 = 10(x+9)(x-3) + 50